OpenStudy (anonymous):

Examine whether  de ned on the set G is a group or not. G = {R}\(1), a * b = a + b + ab, where a,b belongs to R\(1).

terenzreignz (terenzreignz):

Group axioms, you know them? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am having problem with identity

terenzreignz (terenzreignz):

There are four things that must hold for a set and its defined binary operation to be true. A set and binary operation \(\large \left<G,\ast\right>\) is a group if... what?

terenzreignz (terenzreignz):

oh okay. Let's see if it has an identity.

terenzreignz (terenzreignz):

Is there an element e such that \[\Large e\ast a = a = a\ast e\] ?

OpenStudy (anonymous):

its e=0

terenzreignz (terenzreignz):

Then... what's the problem? :D \[\Large 0\ast a = 0 + a + 0a = a = a\ast 0\] ^_^

OpenStudy (anonymous):

when we get e(a+1)=0 either a=-1 or e=0 so when a=-1 then 0/0 comes which is not defined

terenzreignz (terenzreignz):

You want to find \(\large 0 \ast -1\) ? Then... \[\Large 0\ast -1 = 0 +(-1) + (-1\times 0)= -1\] So what's the problem? :)

terenzreignz (terenzreignz):

It just so happened that by definition, \(\large -1 \ast a= -1\) for all a in G. It doesn't have to be a problem :)

OpenStudy (anonymous):

ok thank you